At the beginning of every football game, the referee flips a coin to see who will kick off. The equation that gives the height (in feet) of the coin tossed in the air is h=6+32t−16t^2. To find the times at which the coin is 18 feet above the ground we substitute 18 for h in the equation, giving us
18=6+32t−16t^2


Solve the equation for t to find the times at which the coin is 18 feet above the ground.
The coin is 18 feet above the ground after___________ seconds.

Answer by ewatrrr(24785)

18=6+32t−16t^2
16t^2 -32t + 12 = 0
4t^2 -8t + 3 = 0
(2t - 1)(2t - 3) = 0,
The coin is 18 feet above the ground after 1/2 seconds.